OpenStudy (anonymous):

A 8.0-kg block is sliding down a wall with constant velocity. The coefficient of static friction between the block and the wall is μs = 0.70. The coefficient of kinetic friction between the block and the wall is μk = 0.40. What is the magnitude of the force (in N) pressing the block against the wall? Use g = 9.79 m/s2.

OpenStudy (abhisar):

So, in this case the gravitational pull downwards on the box should be equal to the frictional force exerted on the box by the wall.

OpenStudy (abhisar):



OpenStudy (abhisar):

Since, the block is moving with a constant velocity we will use coefficient of kinetic friction. \(\large \boxed{\tt Frictional~force=Normal~force\times \mu}\) Normal force here is the force by which the box is pressed against the wall. Finally we can make an equation, \(\sf N\times \mu_k=m\times g\) Plug in the values and find the value of N.